# LowDiff: Efficient Frequent Checkpointing via Low-Cost Differential for High-Performance Distributed Training Systems

### v1

@@ III-C Main idea
Findings 1 and 2 show that the compressed gradient can construct the compressed differentials and its size is much smaller than the compressed differential. Thus, the compressed gradient can be leveraged to act as the compressed differential checkpoint of DC, so as to 1) remove the compression stalls via eliminating DC’s differential compression, and 2) reduce or even remove the transmission stalls since the small-size compressed gradient is likely to enable DC’s transmission to be overlapped with the training process.
@@
III-D Data dependency analysis for LowDiff
Requirement of LowDiff. LowDiff requires that gradient reuse be decoupled from training, allowing DC to run in parallel. Without this decoupling, gradient reuse may introduce additional stalls in training (i.e., data dependency [53]). Specifically, in standard training, checkpointing can run independently of the forward and backward passes. However, checkpointing is data dependent on the model update pass [53], since it must read the parameter after model updates (i.e., Write-after-Read (WAR) [21]), which incurs a training stall for synchronization. Prior work [36, 15] also takes the data dependencies into account to design a checkpointing strategy that overlaps with training and reduces stalls. Consequently, it is crucial to carefully examine the coupling introduced by gradients between training and checkpointing. Next, we first show the data dependency when deploying DC directly, and then explore how reusing compressed gradients for DC interacts with the training process.
Deploying DC directly introduces additional data dependencies. As shown in Figure 3(a), checkpointing depends not only on the latest model state but also on the model state from $n$ iterations earlier. This introduces two major issues: 1) The model update in iteration $t+1$ must wait for checkpointing in iteration $t$, leading to a training stall due to the transmission of checkpointing. 2) The model state from iteration tnt-n must be retained in GPU memory until the differential computation in iteration $t$, resulting in extra GPU memory overhead. Moreover, both checkpoint compression and writing lie on the critical path of these data dependencies, and therefore cannot be effectively overlapped with training.
Reusing compressed gradients for DC eliminates such data dependency. Gradients are generated during the backward pass and remain unchanged thereafter. The update pass only reads these gradients (read-only) to update the model state, which removes the Write-after-Read (WAR) dependency and allows gradient reuse to be performed immediately after backpropagation without waiting for the model update. Moreover, by treating the compressed gradient as the compressed differential, the construction of the differential no longer depends on the model states from iterations $t$ or tnt-n, eliminating prior data dependencies for DC calculation. As illustrated in Figure 3(b), together these factors ensure that gradient reuse introduces no data dependency from checkpointing to training, thereby decoupling the two processes and allowing them to execute fully in parallel.
Figure 3: Data dependency analysis.
IV Design
@@
  • Lowering the costs of DC. LowDiff reduces or even eliminates the computation and transmission costs of compressed differential checkpoints by reusing compressed gradients (§IV-A).
  • Enhancing DC performance. LowDiff accelerates the differential checkpointing under high frequency via batched gradient writing optimization (§IV-B).
  • Minimizing the wasted time. LowDiff minimizes the wasted time by analytically tuning the full checkpointing frequency and gradient batching size (§IV-C).
@@
IV-A Reusing Compressed Gradients
@@
Checkpointing process. The checkpointing process repeatedly retrieves the compressed gradients from Reusing Queue $Q$ to CPU memory (Line 11). Subsequently, the checkpointing process determines whether to merge the compressed gradient into batched differential checkpoints (detailed in §IV-B). After that, it invokes torch.save to persist the compressed differential checkpoints $C^{D}_{t}$ (Lines 12-14) to storage. The model state is also regularly checkpointed as a full checkpoint $C^{F}_{t}$ at the $t$-iteration (Line 15).
@@
Algorithm 1 Reusing in LowDiff
Model state $M_{t}$at iteration $t$, Gradient $G_{i,t}$at worker $i$, Reusing queue $Q$, Full checkpoint $C^{F}_{t}$, Differential checkpoint $C^{D}_{n}$at iteration $n$, Compressor CompComp, Decompressor $Comp^{-1}$. Training Process $t$
[…]
IV-B Batched Gradient Writing Optimization
Limitations of LowDiff. We find that LowDiff in §IV-A still has two limitations: Limitation 1: Many writes due to frequent DC. The high frequency leads to numerous differential checkpoints, which incur significant write overhead. Limitation 2: Increased GPU memory consumption from frequent gradient reuse. The high frequency of gradient reuse results in many compressed gradients being retained in GPU memory, waiting to be written to storage.
Batched Write. To address Limitation 1, we can batch multiple writes of compressed gradients (reused as differential checkpoints) into a single write operation, with the help of the widely-used gradient accumulation technique [30, 22, 2] where the gradients of the same shape and size can be accumulated. Note that the above batching scheme, which benefits DC writes, cannot work for traditional checkpointing that only saves full checkpoints, since if a new full checkpoint comes, the previous full checkpoints become obsolete and can be discarded without the need to accumulate them.
Offloading batching to CPU. To address Limitation 2, we adopt a classic GPU-CPU offloading scheme [42] that offloads the batching operation to the CPU and caches the compressed gradients in CPU memory, thus reducing the GPU memory cost caused by frequent checkpointing. Note that the batching operation can be easily handled by the CPU since the operation mainly involves the addition of compressed gradients.
Design details To solve Limitation 1 & 2, we design a batched gradient writing optimization, which offloads multiple compressed gradients from GPU to CPU memory, and then aggregates them into batched writes to storage. As shown in Figure 5, we specify the batched gradient writing optimization of LowDiff in three steps:
@@ IV-B Batched Gradient Writing Optimization
Figure 5: Batched gradient write optimization
IV-C Optimizing Checkpointing Configuration
Insight into configuring LowDiff. Note that recent studies on optimizing the checkpointing performance only focus on configuring the checkpointing frequency (e.g., [19]). However, we observe an insight that LowDiff’s checkpointing performance will also rely on the batching size, i.e., the amount of the batched gradients every batching, introduced in §IV-B. To verify the insight, we take another measurement analysis with the same setting as that in §III-A. Table I shows the results of the wasted time, which evaluates the checkpointing performance (see §II-B), under different batching sizes and checkpointing frequencies for full checkpoint (called full checkpoint frequency, FCF for short). Note that all the results are normalized by comparing them to the lowest wasted time.
We first see that when FCF is too low ($FCF=10$) or too high ($FCF=100$), the wasted time will be increased, since a low FCF will incur high recovery time while a high FCF will incur high steady state overhead (GPU time), Thus, we need to tune FCF to minimize the wasted time, which has been found by Microsoft [19].
@@ IV-C Optimizing Checkpointing Configuration
V enhancement
@@
Issue of LowDiff without gradient compression. The current design of LowDiff relies on gradient compression to enable efficient gradient reuse. When compression is disabled, we can still enforce a lightweight, reuse-only compression to preserve the bandwidth reduction in Finding 2, but the computational savings of Finding 1 no longer apply. In this case, the differential checkpoints become significantly larger, slowing down their snapshotting and persistence accordingly. As a result, checkpointing competes more directly with training for GPU and I/O resources, making overlap optimization even more critical to mitigate the training stall introduced by gradient reuse, snapshotting, and persistence of DC under such non-compressed scenarios. We next present two key insights on snapshotting and persistence, followed by a detailed description of how LowDiff can be enhanced in scenarios without gradient compression (LowDiff+ for short).
V-A Layer-wise gradient reusing and snapshotting.
A key challenge in the non-compression setting arises during the snapshot stage: unlike compressed gradients, transferring raw gradients without compression from GPU to CPU introduces heavier data movement and increases the likelihood of blocking. To accelerate such large-scale snapshotting, we aim to design a fine-grained strategy that better overlaps gradient transmission with training.
As shown in Figure 6, gradients are produced incrementally in the backward pass, following the reverse layer order during training. Instead of waiting until the entire backward pass finishes to snapshot all gradients, which would incur high latency and risk blocking training, modern frameworks (e.g., DeepSpeed [10], PyTorch DDP [28], and Horovod [46]) already exploit layer-wise gradient availability by initiating communication as soon as each layer’s gradients are produced. Inspired by this execution model, we design a fine-grained reuse strategy: immediately snapshot each layer’s gradient to host memory upon computation and stream it into a differential checkpoint assembler. By pipelining backpropagation, communication, and gradient reuse at the layer granularity, we can effectively hide checkpointing overhead behind ongoing training, thereby maximizing overlap with training and reducing stalls.
@@
V-B CPU-based asynchronous persistence strategy.
@@
The layer-wise strategy addresses the snapshotting issue in LowDiff with compressed gradients. Another challenge lies in the persistence of both full and differential checkpoints. The batching strategy described in §IV-B accumulates small compressed updates on the CPU side before writing them out. It also periodically persists the full model state, which works effectively when update sizes are small and sparse. In the non-compression setting, differential checkpoints are significantly larger, and naively batching them demands a sustained write throughput that exceeds the available storage bandwidth, making it essential to design a more efficient checkpoint merging strategy that better lowers the writing data volume.
@@
To address this, we observe that in the non-compression setting, gradients do not need to be persisted separately. Instead, they can be directly fused with the full model state inside CPU memory, enabling the CPU to maintain an always-up-to-date replica of the GPU model. Once gradients are merged into the CPU-resident model state, persistence only needs to occur from the CPU model to storage. By fusing gradients into the CPU-resident model state instead of persisting them separately, redundant writes of differential checkpoints are avoided, which directly reduces the overall storage volume. The different batching strategies of LowDiff and LowDiff+ are shown in 7. This strategy decouples persistence from GPU training, enabling fully asynchronous checkpointing with minimal interference to ongoing training.
V-C Design of LowDiff+

### v2

@@ III-C Main idea
Findings 1 and 2 show that the compressed gradient can construct the compressed differentials, and its size is much smaller than the compressed differential. Thus, the compressed gradient can be leveraged to act as the compressed differential checkpoint of DC, so as to (1) remove the compression stalls via eliminating DC’s differential compression, and (2) reduce or even remove the transmission stalls since the small-size compressed gradient is likely to enable DC’s transmission to be overlapped with the training process.
@@
While LowDiff provides a low-cost DC way based on gradient reuse, its integration into training systems, particularly the feasibility of concurrent execution with the training process, requires further analysis. We next analyze such concurrency in §IV.
IV Analysis
By incorporating the main idea of gradient reuse (§III), LowDiff aims to enable the DC process to run concurrently with model training, thereby minimizing potential stalls. To confirm the feasibility of such parallelism, we analyze our proposed strategy from two key aspects:
  • Parallelism property: it ensures no data dependencies block concurrency with training iterations.
  • Execution time overlap: it reveals whether the checkpointing cost can be effectively hidden within the training time.
We provide a detailed analysis of these two aspects as follows.
Figure 3: LowDiff’s DC can run in parallel with forward pass(F), backward pass(B), and model update(U).
IV-A Parallelism Analysis
Necessity of Parallelism Analysis. Checkpointing can be executed with the forward and backward passes which do not modify the model state [mohan2021checkfreq], but it cannot run in parallel with model update which directly writes to it [wang2024fastpersist]. In other words, checkpointing is data-dependent on the model update [wang2024fastpersist]. The reason is that it must read the model state after model updates (Read-after-Write [hennessy2011computer], RAW dependency), and the model update at the next iteration must write the model state after checkpointing completes its read (Write-after-Read [hennessy2011computer], WAR dependency). Consequently, the two dependency patterns incur training stalls, preventing from paralleling checkpointing with model update. Prior studies [mohan2021checkfreq, wang2023gemini, wang2024fastpersist] design their checkpointing strategies based on these data dependencies to pipeline training and checkpointing.
Consequently, it is necessary to carefully analyze: (1) the data dependency introduced by our reuse strategy and (2) the parallelism property of LowDiff’s DC, specified as follows.
(1) LowDiff’s DC can eliminate the WAR dependency. We next analyze current DC’s WAR dependency with next iteration’s model update, backward, and forward pass.
  • Model update: as described in §II-A, the model update (Equation (4)) only requires read-only access to the gradient. Because LowDiff checkpoints only the compressed gradient, leveraging gradient reuse, the original Write-after-Read (WAR) dependency is eliminated because the checkpointing process no longer needs to read the model state ($M_{t}$). Instead, it just reads the compressed gradient ($\widetilde{G}_{t}$), which is available in the backward pass and will not be changed after. Since both the checkpointing and model update are read-only consumers of the gradient, they can operate in parallel without conflict.
  • Backward pass: LowDiff’s gradient reuse strategy introduces no additional WAR dependency on backward pass, since each gradient ${G}_{t}$ is finalized within its iteration and never modified thereafter. Specifically, once a gradient ${G}_{t}$ is generated in the current iteration ($t$), it is final and will not be modified by subsequent steps. The backward pass of the $t+1$-th iteration (Equation (2)) does not read or depend on ${G}_{t}$; it generates a completely new gradient ($G_{t+1}$) from scratch and won’t change the ${G}_{t}$. Therefore, the DC’s read of ${G}_{t}$ does not conflict with the $B$’s write of ${G}_{t+1}$ at $t+1$-th iteration, allowing LowDiff’s DC to run in parallel with the next iteration’s backward pass.
  • Forward pass: The forward pass (Equation (1)) doesn’t involve gradient, so there is no WAR dependency between LowDiff’s DC and forward pass.
(2) LowDiff can run in parallel with forward pass, backward pass, and model update. To visualize the parallelism benefit from data dependency, Figure 3 shows the execution timeline with LowDiff. We see that the DC task ($C^{D}_{t}$) at the $t$-th iteration is no longer coupled with the model update ($U$) or the following iteration (no dependency arrow from DC to training). Instead, the reusing operation on compressed gradient ($\widetilde{G}_{t}$) begins immediately after the backward pass ($B$), running in parallel with the model update and subsequent iteration’s forward pass(F), backward pass (B), and model update(U). This parallelism property effectively transforms LowDiff’s DC from a blocking operation to a parallel execution with training, thereby eliminating the synchronization stalls previously caused by the WAR dependency.
Figure 4: Time of iteration, full checkpointing, and differential checkpointing.
IV-B Overlap Analysis
Necessity of Overlap Analysis. Beyond its parallelism property, LowDiff can perform even better if the checkpointing process is fast enough to be hidden by the training iteration, thus minimizing DC’s cost. Specifically, the time required to checkpoint a compressed gradient (denoted as $C$) should be shorter than a single training iteration (denoted as $F+B+U$), similar to the Microsoft study [mohan2021checkfreq].
Consequently, it is also necessary to analyze the potential of overlapping, so we first (1) take a measurement analysis of training time, full checkpointing, and DC time, as well as (2) discuss why the overlap can be practically feasible.
(1) The checkpointing of compressed gradients can be overlapped with training. Following the measurement setup in §III-A, we measure the iteration time, full checkpointing time and differential checkpointing time on BERT-B, BERT-L, GPT2-S, and GPT2-L. As shown in Figure 4, the DC time is only 20.5%, 22.1%, 23.1%, and 24.6% of the corresponding training iteration time for these models. This indicates that in most cases, the checkpointing process can be fully overlapped with training iterations without introducing noticeable stalls.
(2) Modern hardware makes the overlap practically feasible. NVMe SSDs based on PCIe 4.0 offer write speeds around 5 GB/s (and up to 12 GB/s for PCIe 5.0, with RAID 0 configurations being even faster). Given that LowDiff’s compressed-gradient checkpoints are relatively small (several hundred MBs to a few GBs, as evaluated in §VIII), the theoretical I/O write time is typically under 1 second. This duration is significantly shorter than that of a full training iteration [wang2023gemini].
Based on the above analysis, we confirm that the concurrent execution between LowDiff’s DC and training is theoretically feasible. While this analysis establishes the feasibility of frequent DC, the detailed design of LowDiff will be presented in §V.
V Design
@@
  • Lowering the costs of DC. LowDiff reduces or even eliminates the computation and transmission costs of compressed differential checkpoints by reusing compressed gradients (§V-A).
  • Enhancing DC performance. LowDiff accelerates the differential checkpointing under high frequency via batched gradient writing optimization (§V-B).
  • Minimizing the wasted time. LowDiff minimizes the wasted time by analytically tuning the full checkpointing frequency and gradient batching size (§V-C).
@@
V-A Reusing Compressed Gradients
@@
Algorithm 1 Reusing in LowDiff
1: Model state $M_{t}$ at iteration $t$, Gradient $G_{i,t}$ at worker $i$, Reusing queue $Q$, Full checkpoint $C^{F}_{t}$, Differential checkpoint $C^{D}_{n}$ at iteration $n$, Compressor CompComp, Decompressor $Comp^{-1}$.
2: function Training Process
3:   for all $t$ do
4:    $G_{i,t}\leftarrow Backward(x_{t})$
5:    $\widetilde{G}_{i,t}\leftarrow Comp(G_{i,t})$ $\triangleright$ Compress
6:    $\widetilde{G}_{t}\leftarrow Sync(\widetilde{G}_{i,t})$ $\triangleright$ Synchronize
7:    ${\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.put}}(\widetilde{G}_{t})$ $\triangleright$ Zero-Copy
8:    $G_{t}\leftarrow Comp^{-1}(\widetilde{G}_{t})$ $\triangleright$ Decompress
9:    $M_{t+1}\leftarrow M_{t}+Adam(o_{t},G_{t})$ $\triangleright$ Update
10: function Checkpointing Process
11:   while $\widetilde{G}_{t}\leftarrow{\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.get}}$ do $\triangleright$ Reusing
12:    $C^{D}_{t}\leftarrow Save(\widetilde{G}_{t})$ $\triangleright$ Diff CKPT
13:   $C^{F}_{t}\leftarrow Save(M_{t})$ $\triangleright$ Full CKPT Regularly
14: function Recovery Process
15:   $M_{t}\leftarrow Load(C^{F}_{t})$ $\triangleright$ Load Full CKPT
16:   for $j\leftarrow t$, …, $n$ do $\triangleright$ Restore to Latest
17:    $\widetilde{G}_{j}\leftarrow Load(C^{D}_{j})$ $\triangleright$ Load Diff CKPT
18:    $G_{j}\leftarrow Comp^{-1}(\widetilde{G}_{j})$ $\triangleright$ Decompress
19:    $M_{j+1}\leftarrow M_{j}+Adam(G_{j})$ $\triangleright$ Diff Merge
20:   return $M_{n}$
@@
Checkpointing process. The checkpointing process repeatedly retrieves the compressed gradients from Reusing Queue $Q$ to CPU memory (Line 11). Subsequently, the checkpointing process determines whether to merge the compressed gradient into batched differential checkpoints (detailed in §V-B). After that, it invokes torch.save to persist the compressed differential checkpoints $C^{D}_{t}$ (Lines 12-14) to storage. The model state is also regularly checkpointed as a full checkpoint $C^{F}_{t}$ at the $t$-iteration (Line 15).
@@
V-B Batched Gradient Writing Optimization
Limitations of LowDiff. We find that LowDiff in §V-A still has two limitations:
  • Limitation 1: Many writes due to frequent DC. The high frequency leads to numerous differential checkpoints, which incur significant write overhead.
  • Limitation 2: As analyzed in §IV-A, avoiding the WAR dependency requires temporarily retaining the compressed gradient $\widetilde{G}_{t}$ for the DC operation. Consequently, frequent DCs cause this gradient buffer to remain occupied while waiting for the write to storage, leading to extra GPU memory consumption.
we provide the batched write and offloading strategy to address the corresponding limitation as follows.
Batched write. To address Limitation 1, we can batch multiple writes of compressed gradients (reused as differential checkpoints) into a single write operation, with the help of the widely-used gradient accumulation technique [lin2017deep, hermans17a, accumulation] where the gradients of the same shape and size can be accumulated. Note that the above batching scheme, which benefits DC writes, cannot work for traditional checkpointing that only saves full checkpoints, since if a new full checkpoint comes, the previous full checkpoints become obsolete and can be discarded without the need to accumulate them.
Offloading batching to CPU. To address Limitation 2, we adopt a classic GPU-CPU offloading scheme [rajbhandari2020zero] that offloads the batching operation to the CPU and caches the compressed gradients in CPU memory, thus reducing the GPU memory cost caused by frequent checkpointing. Note that the batching operation can be easily handled by the CPU since the operation mainly involves the addition of compressed gradients.
Figure 6: Batched gradient write optimization
Design details. To solve Limitation 1 & 2, we design a batched gradient writing optimization, which offloads multiple compressed gradients from GPU to CPU memory, and then aggregates them into batched writes to storage. As shown in Figure 6, we specify the batched gradient writing optimization of LowDiff in three steps:
@@ V-B Batched Gradient Writing Optimization
V-C Optimizing Checkpointing Configuration
Insight into configuring LowDiff. Note that recent studies on optimizing the checkpointing performance only focus on configuring the checkpointing frequency (e.g., [gupta2024just]). However, we observe an insight that LowDiff’s checkpointing performance will also rely on the batching size, i.e., the amount of the batched gradients every batching, introduced in §V-B. To verify the insight, we take another measurement analysis with the same setting as that in §III-A. Table I shows the results of the wasted time, which evaluates the checkpointing performance (see §II-B), under different batching sizes and checkpointing frequencies for full checkpoint (called full checkpoint frequency, FCF for short). Note that all the results are normalized by comparing them to the lowest wasted time.
We first see that when FCF is too low ($FCF=10$) or too high ($FCF=100$), the wasted time will be increased, since a low FCF will incur high recovery time while a high FCF will incur high steady state overhead (GPU time), Thus, we need to tune FCF to minimize the wasted time, which has been found by Microsoft [gupta2024just].
@@ V-C Optimizing Checkpointing Configuration
VI enhancement
The current design of LowDiff relies on gradient compression to enable efficient gradient reuse. However, in some high-precision training scenarios, any potential accuracy degradation caused by compression is unacceptable. If the distributed training system does not inherently employ a gradient compression strategy, making it impossible to leverage compression-induced bandwidth advantages. So we also need to design a non-compression high-frequency DC strategy for above scenarios. We next present two key insights on snapshotting and persistence, followed by a detailed description of how LowDiff can be enhanced in scenarios without gradient compression (LowDiff+ for short).
@@
VI-A Layer-Wise Gradient Reusing and Snapshotting Strategy.
Issue of snapshotting. A key issue in the non-compression setting arises during the snapshot stage: unlike compressed gradients, transferring raw gradients without compression from GPU to CPU introduces heavier data movement and increases the likelihood of a stall. The snapshotting competes more directly with training for GPU and I/O resources, making overlap optimization even more critical to mitigate the training stall. To accelerate such large-scale snapshotting, we aim to design a fine-grained strategy that better overlaps gradient transmission with training.
Insight 1: gradient can be reused layer-by-layer. We recall that gradients are produced incrementally in the backward pass, following the reverse layer order during training, as shown in Figure 7. Instead of waiting until the entire backward pass finishes to synchronize all gradients, which would incur high latency and risk blocking training, modern frameworks (e.g., DeepSpeed [deepspeed], PyTorch DDP [li2020pytorch], and Horovod [sergeev2018horovod]) already exploit layer-wise gradient availability by initiating communication as soon as each layer’s gradients are produced. Inspired by this execution model, we design a fine-grained reuse strategy: once a layer’s gradient has been synchronized, we immediately reuse it to the reusing queue, snapshot it to host memory, and stream it into a differential checkpoint assembler. This granular approach enables parallel data movement, allowing concurrent snapshot writes to CPU DRAM to better utilize memory bandwidth, thereby mitigating the high transfer volume. By pipelining backpropagation, communication, gradient reuse, and snapshotting at the layer granularity, we can effectively hide checkpointing overhead behind ongoing training, thereby maximizing overlap with training and reducing stalls.
@@
VI-B CPU-Based Asynchronous Persistence Strategy.
@@
Issue of persistence. While the layer-wise strategy addresses the snapshotting issue in the non-compression setting, another issue remains: persistence of both full and differential checkpoints. The batching strategy described in §V-B accumulates small compressed updates on the CPU side before writing them out. It also periodically persists the full model state, which works effectively when update sizes are small and sparse. In the non-compression setting, differential checkpoints are significantly larger, and naively batching them demands a sustained write throughput that exceeds the available storage bandwidth, making it essential to design a more efficient checkpoint merging strategy that better lowers the writing data volume.
@@
Insight 2: differential and full checkpoints can be batched together in CPU memory. We find that in the non-compression setting, the differential checkpoints (i.e. gradients) and full checkpoints do not need to be persisted separately. Instead, they can be directly fused with the full model state inside CPU memory, enabling the CPU to maintain an always-up-to-date replica of the GPU model. Once gradients are merged into the CPU-resident model state, persistence only needs to occur from the CPU model to storage. By fusing gradients into the CPU-resident model state instead of persisting them separately, redundant writes of differential checkpoints are avoided, which directly reduces the overall storage volume. The different batching strategies of LowDiff and LowDiff+ are shown in Figure 8. This strategy decouples persistence from GPU training, enabling fully asynchronous checkpointing with minimal interference to ongoing training.
VI-C Design of LowDiff+